OpenStudy (anonymous):

can someone help? differtiate4/[ln(2x^3-5)^3

OpenStudy (anonymous):

that the sound of the men.. working on the chain..

OpenStudy (dumbcow):

Yes use the chain rule. Try to really break it down into more manageable functions: Say \[w = \frac{4}{u^{3}} \rightarrow dw = -\frac{12}{u^{4}}\] \[u = \ln v \rightarrow du = \frac{1}{v}\] \[v = 2x^{3}-5 \rightarrow dv = 6x^{2}\] Then put it all together in terms of x \[f'(x) = dw*du*dv = -\frac{72x^{2}}{v*u^{4}} = -\frac{72x^{2}}{(2x^{3}-5)*\ln^{4}(2x^{3}-5)}\]